OpenStudy (lizz123):

Medal @freckles Perform the operations using adding and subtracting polynomials.

OpenStudy (lizz123):

\[7x - [2\left( x ^{2} -z\right) + 4x ^{2} -7z] + 6z ^{2}\]

OpenStudy (freckles):

ok so let's do any distributing needed a(b+c)=ab+ac

OpenStudy (freckles):

let's begin with the 2(x^2-z) thing

OpenStudy (freckles):

can apply the distributive property there

OpenStudy (freckles):

\[2(x^2-z) \\ \text{ here we are just going \to multiply the two terms inside the ( ) \by 2 } \\ 2(x^2)-2(z)=2x^2-2z\]

OpenStudy (freckles):

so now you have \[7x-[2x^2-2z+4x^2-7z]+6z^2\]

OpenStudy (freckles):

you have one more area to apply distributive property

OpenStudy (freckles):

do you see the - in front of the [ ]?

OpenStudy (freckles):

drop the [ ] and just change the sign to each term inside the [ ] this is the same as multiplying by -1

OpenStudy (freckles):

I will let you try this part

OpenStudy (lizz123):

So, \[7x - [ (2x^{2} -2z)+4x ^{2} -7z] +6\]

OpenStudy (lizz123):

\[7x + [ -2x ^{2}+2z -4x ^{2}+7z]+6\]

OpenStudy (freckles):

yep

OpenStudy (freckles):

now combine like terms

OpenStudy (freckles):

\[+7x\\ -2x^2-4x^2=? \\ 2z+7z=? \\+ 6\]

OpenStudy (lizz123):